### v1

# On hyperbolic links associated to Eulerian cycles on ideal right-angled hyperbolic $3$-polytopes

Abstract.
We consider Eulerian cycles without transversal selfintersections in $4$-valent planar graphs. We prove that any cycle of this type in the graph of an ideal right-angled hyperbolic $3$-polytope corresponds to a hyperbolic link such that its complement consists of $4$-copies of this polytope glued according to its checkerboard coloring. Moreover, this link consists of trivially embedded circles bijectively corresponding to the vertices of the polytope. For such cycles we prove that the $3$-antiprism $A(3)$ (octahedron) has exactly $2$ combinatorially different cycles, the $4$-antiprism $A(4)$ has exactly $7$ combinatorially different cycles, and these cycles correspond to $7$ cycles (perhaps combinatorially equivalent) on any polytope different from antiprisms, and any antiprism $A(k)$ has at least $2$ combinatorially different cycles. The $2$-fold branched covering space corresponding to our link is a small cover over some simple $3$-polytope. This small cover is defined by a Hamiltonian cycle on it. We show that any compact hyperbolic $3$-polytope with a Hamiltonian cycle arises in this way.
Key words and phrases:
convex polytope, link, chain, ideal right-angled hyperbolic polytope, small cover
2010 Mathematics Subject Classification:
57S12, 57S17, 57S25, 52B05, 57K10, 52B70, 57K10
1. Introduction
The theory of knot and links is a classical area of mathematics developing since XIX century. One of the well-known directions in this area is the theory of hyperbolic links. These are links with the complement having the structure of a complete hyperbolic manifold. In this paper using methods and results of toric topology we build a wide family of hyperbolic links corresponding to Eulerian cycles on any ideal right-angled hyperbolic $3$-polytope. Usually ideal right-angled polytopes arise when the alternating diagram of a link is reduced to some canonical form (see, for example [CKP21]). In our approach we build a link with trivial circles corresponding to vertices of the ideal right-angled polytope and their structure is defined by the Eulerian cycle. Any $k$-antiprism has a canonical Eulerian cycle. In this case our decomposition of the complement to the $(2k)$-link chain into $4$ antiprisms coincides with the decomposition described by W.P. Thurston [T02, Example 6.8.7].
Our main result (Theorem 3.2) is that any Eulerian cycle without transversal selfintersections in the graph of any ideal right-angled $3$-polytope $P$ corresponds to a link whose circles bijectively correspond to vertices of $P$ and the complement is decomposed into $4$ copies of $P$. We show that the $2$-fold branched covering space corresponding to this link is a small cover build by the A.D. Mednykh [M90] construction from a Hamiltonian cycle on a simple $3$-polytope. We provide a method how to build such cycles on any ideal right-angled $3$-polytope.
2. Basic facts
Definition 2.1.
A cycle in a graph $G$ is called Eulerian if it passes each edge of the graph once (it may pass one vertex many times). We call an Eulerian cycle in the $4$-valent planar graph nonselfcrossing if at each vertex it does not intersect itself transversally, that is each time it visits the vertex it turns left or right, but does not go straight.
A cycle in a graph is called Hamiltonian, if it visits each vertex once.
Let $G$ be a planar. Its medial graph is a new planar graph $M(G)$ with vertices bijectively corresponding to edges of $G$. Its edges arise when we walk around the boundary cycle of each face. Each vertex of this cycle corresponds to an edge of $M(G)$ connecting the vertices corresponding to successive edges of the cycle. A medial graph is $4$-valent. Is is known that a graph $G$ is a graph $G(P)$ of an ideal hyperbolic right-angled $3$-polytope $P$ if and only if it is a medial graph of some simple polytope $\widehat{P}$. Moreover, $\widehat{P}$ is defined uniquely up to passing to the dual polytope $\widehat{P}^{*}$ (see more details in [E19]).
Proposition 2.2.
Nonselfcrossing Eulerian cycles in a $4$-valent planar graph $G$ correspond to Hamiltonian cycles in $M(G)$.
Each ideal right-angled $3$-polytope $P$ admits a checkerboard coloring: its faces can be colored in black and white colors in such a way that adjacent faces have different colors (if $)G(P)=M(G(\widehat{P}$)), then black faces of $P$ correspond to vertices of $\widehat{P}$ and white facets of $P$ correspond to facets of $\widehat{P}$). Assign to white color the vector $e_{1}\in\mathbb{Z}_{2}^{2}$ and to black color $e_{2}\in\mathbb{Z}_{2}^{2}$. Then we obtain the mapping $\Lambda_{P}\colon\{F_{1},\dots,F_{m}\}\to\mathbb{Z}_{2}^{2}$, $F_{i}\to\Lambda_{i}$, from the set of facets of $P$ to $\mathbb{Z}_{2}^{2}$, and A.Yu.Vesnin–A.D. Mednykh construction (see [V17]) gives the complete hyperbolic manifold $N(P)$ of finite volume glued of $4$ copies of $P$:
$N(P)=P\times\mathbb{Z}_{2}^{2}/\sim,(p,a)\sim(q,b)\text{ if and only if }p=q\text{ and }a-b\in\langle\Lambda_{i}\colon p\in F_{i}\rangle.$
In this formula the ideal vertices are not assumed to belong to $P$.
Definition 2.3.
A hyperelliptic manifold $M^{n}$ is an $n$-manifold with an action of an involution $\tau$ such that $M^{n}/\langle\tau\rangle$ is homeomorphic to $S^{n}$. The involution $\tau$ is called hyperelliptic.
In [M90] A.D. Mednykh constructed examples of hyperelliptic $3$-manifolds with geometric structures modelled on five of eight Thurston’s geometries: $\mathbb{R}^{3}$, $\mathbb{L}^{3}$, $\mathbb{S}^{3}$, $\mathbb{L}^{2}\times\mathbb{R}$, and $\mathbb{S}^{2}\times\mathbb{R}$. Each example was built using a right-angled $3$-polytope $P$ equipped with a Hamiltonian cycle. This construction can be described as follows.
Let $\Gamma$ be a Hamiltonian cycle in the graph of a simple $3$-polytope $Q$. Then it divides $\partial Q\simeq S^{2}$ into two disks. Each edge of $Q$ not lying in $\Gamma$ divides one of the disks into two disks. Thus, the adjacency graph of faces of $Q$ lying in the closure of each component of $\partial Q\setminus\Gamma$ is a tree and these faces can be colored in two colors in such a way that adjacent faces have different colors. Then combining both components we obtain a coloring of faces of $Q$ in four colors. This coloring corresponds to a mapping $\widetilde{\Lambda}_{\Gamma}$ of the set of faces of $Q$ to $\mathbb{Z}_{2}^{3}$ by the rule: the first three colors correspond to basic vectors $e_{1},e_{2},e_{3}\in\mathbb{Z}_{2}^{3}$, and the fourth corresponds to their sum $e_{1}+e_{2}+e_{3}$. This mapping gives an orientable $3$-manifold $N(Q,\widetilde{\Lambda}_{\Gamma})$ (it toric topology [BP15] it is called an orientable small cover):
$N(Q,\widetilde{\Lambda}_{\Gamma})=Q\times\mathbb{Z}_{2}^{3}/\sim,(p,a)\sim(q,b)\text{ if and only if }p=q\text{ and }a-b\in\langle\Lambda_{i}\colon p\in F_{i}\rangle.$
If the components of $\partial Q\setminus\Gamma$ correspond to the pairs of colors $(1,2)$ and $(3,4)$, then $\tau=e_{1}+e_{2}=e_{3}+(e_{1}+e_{2}+e_{3})$ is the hyperelliptic involution on $N(Q,\widetilde{\Lambda}_{\Gamma})$. Moreover, the mapping $N(P,\widetilde{\Lambda}_{\Gamma})\to S^{3}$ is a $2$-fold branched covering with the following branch set (see details in [EE25, Section 4.5]). The edges of $Q$ not lying in $\Gamma$ form a perfect matching of the graph $G(Q)$ – a disjoint set of edges covering all the vertices. Then the preimage of this set in $N(Q,\widetilde{\Lambda}_{\Gamma})$ and in $S^{3}$ is a disjoint set of circles, each circle glued of two copies of the corresponding edge. This link is the brach set of the covering. The detailed description of this link is given in [G24]. In [VM99S2] this construction was generalized from Hamiltonian cycles to Hamiltonian theta-graphs and Hamiltonian $K_{4}$-subgraphs.
On the language of toric topology this construction and its generalizations is described in [E24] and [EE25]. In particular, in these papers the space corresponding to general vector-coloring of rank $r$ is considered.
Definition 2.4.
A vector-coloring of rank $r$ of a simple $3$-polytope $Q$ is a mapping $\Lambda$ from the set of its facets $F_{1},\dots,F_{m}$ to $\mathbb{Z}_{2}^{r}$, $F_{i}\to\Lambda_{i}$, such that $\langle\Lambda_{1},\dots,\Lambda_{m}\rangle=\mathbb{Z}_{2}^{r}$. It corresponds to the space
$N(Q,\Lambda)=Q\times\mathbb{Z}_{2}^{r}/\sim,(p,a)\sim(q,b)\text{ if and only if }p=q\text{ and }a-b\in\langle\Lambda_{i}\colon p\in F_{i}\rangle.$
This space has an action of $\mathbb{Z}_{2}^{r}$. This space is a manifold if and only if for each vertex $v=F_{i}\cap F_{j}\cap F_{k}$ either $\Lambda_{i}=\Lambda_{j}=\Lambda_{k}$, or two of these vectors are equal and the third is different, or three vectors are linearly independent (see [E24]). In particular, any simple cycle $\Gamma$ in the graph of $Q$ divides $\partial Q$ in two connected components. Define $\Lambda_{i}=e_{1}$, if $F_{i}$ lies in the closure of the first component, and $\Lambda_{i}=e_{2}$, if $F_{i}$ lies in the closure of the second component. This vector-coloring $\Lambda_{\Gamma}$ has rank $2$ and $N(Q,\Lambda_{\Gamma})\simeq S^{3}$ glued of $4$ copies of $Q$ in the following way. $\partial Q$ is divided by $\Gamma$ into two facets. The complex on the polytope $Q$ with these two facets is homeomorphic to the $3$-ball with the boundary divided into two hemispheres by the equator. Then the ball $Q\times(0,0)$ is glued to $Q\times(1,0)$ along the hemisphere corresponding to $e_{1}$ to give a new ball, as well as $Q\times(0,1)$ to $Q\times(1,1)$, while the resulting balls are glued along the boundaries to give the $3$-sphere.
3. Main construction
Construction 3.1.
Let $\gamma$ be a nonselfcrossing Eulerian cycle on the $4$-valent 33-polytope $P$. Let us build a simple $3$-polytope $Q(P,\gamma)$ with a Hamiltonian cycle $\Gamma_{\gamma}$ by the following rule. Substitute each vertex of the graph of $P$ by two vertices connected by an edge in such a way that each pair of successive edges of $\gamma$ at this vertex is incident to the same vertex of the new edge. The new graph satisfies the condition that each face is bounded by a simple edge-cycle, and if two boundary cycles of faces intersect, then by an edge. Thus by the Steinitz theorem this graph is a graph of a unique combinatorial simple polytope $Q(P,\gamma)$ (see more details in [E19]). Moreover, $\gamma$ corresponds to a Hamiltonian cycle $\Gamma_{\gamma}$ on this polytope. The new edges form a perfect matching in the graph of $Q$.
Theorem 3.2.
Let $P$ be an ideal right-angled hyperbolic $3$-polytope. Each nonselfcrossing Eulerian cycle $\gamma$ corresponds to a link $C_{\gamma}\subset S^{3}$ consisting of $\{\text{vertices of }P\}$ circles such that its complement $S^{3}\setminus C_{\gamma}$ is homeomorphic to the hyperbolic manifold $N(P)$ glued of $4$ copies of $P$. Moreover, $C_{\gamma}$ is the branch set of the $2$-fold brach covering $N(Q(P,\gamma),\widetilde{\Lambda}_{\Gamma_{\gamma}})\to S^{3}$.
Proof.
Indeed, the $2$-fold branched covering has the form $N(Q,\widetilde{\Lambda}_{\Gamma_{\gamma}})\to N(Q,\Lambda_{\Gamma_{Q}})$: $p,a]\to[p,\pi(a)]$, where $\pi\colon\mathbb{Z}_{2}^{3}\to\mathbb{Z}_{2}^{3}/\langle e_{1}+e_{2}\rangle\simeq\mathbb{Z}_{2}^{2}$ is a projection. Under the homeomorphism $N(Q,\Lambda_{\Gamma_{Q}})\simeq S^{3}$ the preimage of the perfect matching on $Q$ corresponds to the link $C_{\gamma}$ consisting of $\{\text{vertices of }P\}$ circles (each circle corresponds to an edge of the perfect matching and to a vertex of $P$). $Q\setminus\{\text{perfect matching}\}$ is homeomorphic to $P$, while $S^{3}\setminus C_{\gamma}$ is homeomorphic to $N(P)$. ∎
Remark 3.3.
A link $L$ is called hyperbolic if its complement $S^{3}\setminus L$ has a structure of a complete hyperbolic manifold. In [CKP21] a hyperbolic link was called right-angled, if $S^{3}\setminus L$ with the complete hyperbolic structure admits a decomposition into ideal hyperbolic right-angled polytopes. By construction the link $C_{\gamma}$ is right-angled.
Remark 3.4.
For nonselfcrossing Eulerian cycles on any $4$-valent convex $3$-polytope the analog of Theorem 3.2 without hyperbolic structure on $S^{3}\setminus C_{\gamma}$ holds.
Example 3.5.
The octahedron is a unique right-angled polytope with the smallest number of vertices (equal to $6$). Up to combinatorial symmetries it has exactly two nonselfcrossing Eulerian cycles (see the proof in Fig. 1) shown in Fig. 2. We also present the corresponding simple polytopes and hyperbolic links. These links are not isotopic, but their complements are homeomorphic. The ways to represent links follows the method from [T25].
Figure 1. Enumeration of nonselfcrossing Eulerian cycles on the octahedron
Figure 2. Hyperbolic links corresponding to nonselfcrossing Eulerian cycles on the octahedron
Example 3.6.
Example 3.5 can be generalized as follows. It is known that any antiprism $A(k)$ ($A(3)$ is the octahedron) is an ideal right-angled $3$-polytope (see [V17]). In Fig 3 we show two different nonselfcrossing Eulerian cycles on this polytope. The hyperbolic structure on the complement to the link corresponding to the left cycle is exactly the structure defined by W.P.Thurston in [T02, Example 6.8.7]. We learn this example due to the lectures by A.Yu.Vesnin and the diploma works by D.V. Chepakova [C23] and D.A. Tsygankov [T25]. The case of $A(3)$ is also mentioned in [V17, Section 5.1].
Figure 3. Nonselfcrossing Eulerian cycles on the antiprism
Example 3.7.
It can be shown (see Fig. 4) that up to combinatorial symmetries $A(4)$ has exactly 77 nonselfcrossing Eulerian cycles shown in Fig. 5.
Figure 4. Enumeration of nonselfcrossing Eulerian cycles on $A(4)$
Figure 5. Nonselfcrossing Eulerian cycles on $A(4)$ and the corresponding polytopes
4. Edge-twists and the existence of nonselfcrossing Eulerian cycles
In [V17, Theorem 2.14] on the base of results from [BGGMTW05] the following construction of all the ideal right-angled polytopes was described .
Definition 4.1.
An operation of an edge-twist is shown in Fig. 6. Two edges on the left belong to one facet of a the ideal right-angled polytope and connect $4$ distinct vertices. The result is again an ideal right-angled polytope. Let us call an edge-twist restricted if both edges are adjacent to the same edge, that is the $4$ vertices follow each other during the round walk along the boundary of a facet.
Figure 6. An operation of an edge-twist
Theorem 4.2 ([V17]).
Any ideal right-angled $3$-polytope can be obtained by operations of an edge-twist from some $k$-antiprism $A(k)$, $k\geqslant 3$.
Remark 4.3.
Operations of an edge-twist are not applicable to the octahedron, hence all the other polytopes are obtained from $k$-antiprisms, $k\geqslant 4$.
In [E19, Theorem 9.13] this result was improved.
Theorem 4.4 ([E19]).
A $3$-polytope is an ideal right-angled $3$-polytope if and only if either it is a $k$-antiprism $A(k)$, $k\geqslant 3$, or it can be obtained from the $4$-antiprism by operations of a restricted edge-twist.
The following result is straightforward from the definitions.
Proposition 4.5.
Any edge-twist transforms a nonselfcrossing Eulerian cycle to a nonselfcrossing Eulerian cycle on the new polytope.
Corollary 4.6.
The $3$-antiprism $A(3)$ (octahedron) has exactly $2$ combinatorially different nonselfcrossing Eulerian cycles, the $4$-antiprism $A(4)$ has exactly $7$ combinatorially different nonselfcrossing Eulerian cycles, and they correspond to $7$ nonselfcrossing Eulerian cycles (perhaps some of them are combinatorially equivalent) on any polytope different from antiprisms, and any antiprism $A(k)$ has at least $2$ combinatorially different cycles.
Proof.
This follows from Examples 3.5, 3.6, and 3.7. ∎
Remark 4.7.
As was mentioned by A.A. Gaifullin, the existence of a nonselfcrossing Eulerian cycle can be proved as follows. Since each vertex of $P$ has even valency, it has a Eulerian cycle. Then we can deform this cycle at bad vertices. If it has a transversal self-crossing, then one of the two ways to change this crossing leaves the cycle connected.
Problem 1.
To enumerate all combinatorially different nonselfcrossing Eulerian cycles on any ideal right-angled $3$-polytope. To find estimates for their number.
5. Links associated to Hamiltonian cycles on compact right-angled $3$-polytopes
Proposition 5.1.
Every compact right-angled $3$-polytope $Q$ with a Hamiltonian cycle $\Gamma$ defines an ideal right-angled polytope $P$ obtained by shrinking to points the edges of the perfect matching consisting of the edges not in the cycle. The polytope $Q$ has an induced nonselfcrossing Eulerian cycle $\gamma$ such that $\Gamma=\Gamma_{\gamma}$. For the link $C_{\gamma}$ both the complement $S^{3}\setminus C_{\gamma}$ and the $2$-fold branched covering space have complete hyperbolic structures obtained by gluing right-angled polytopes.
Proof.
Indeed, by [E19, Theorem 11.6] the polytope obtained by cutting off all these edges is almost Pogorelov, and by [E19, Theorem 6.5] shrinking the obtained quadrangles to points give the ideal right-angled $3$-polytope. ∎
Example 5.2.
The Dodecahedron is a unique compact right-angled polytope with minimal number of facets (equal to $12$). Up to combinatorial symmetries it has a unique Hamiltonian cycle. In Fig. 7 we show this Hamiltonian cycle, and the way how the associated ideal right-angled polytope is obtained from $A(4)$ by a sequence of two restricted edge-twists. We also show another polytope corresponding to another Eulerian cycle on the $4$-antiprism. The corresponding links have homeomorphic complements, but the first link has $2$-fold branch covering space with hyperbolic structure, and the $2$-fold branch covering space contains incompressible tori corresponding to $4$-belts – cyclic sequences of $4$ facets such that two facets are adjacent if and only if they are successive (see more details in [E22]).
Figure 7. Hamiltonian cycle on the dodecahedron
Example 5.3.
A fullerene is a simple $3$-polytope with only pentagonal and hexagonal faces. It is known that any fullerene is a right-angled hyperbolic polytope and the dodecahedron is the fullerene with minimal number of facets (see more details in [E19]). It was shown by F. Kardoš [K20] that any fullerene has a Hamiltonian cycle. Each Hamiltonian cycle on a fullerene corresponds to a hyperbolic link with hyperbolic $2$-fold brach covering space.
Problem 2.
To characterise ideal right-angled $3$-polytopes corresponding to Hamiltonian cycles on compact right-angled $3$-polytopes.
Remark 5.4.
For Hamiltonian cycles on general simple $3$-polytopes the analog of Proposition 5.1 is not valid. If we shrink all the edges of the perfect matching complementary to the cycle to points, the resulting graph may not represent a polytope. For example, if $Q$ contains a triangle, then the resulting spherical complex contains a bigon. For a Hamiltonian cycle on the cube the resulting complex also contains a bigon. But if the resulting spherical graph contains no bigonal faces (for example when $Q$ has no triangles and quadrangles), then by the Steinitz theorem it is the graph of a $4$-valent convex polytope. But this polytope may be not ideal right-angled.
Problem 3.
To characterise Hamiltonian cycles on simple $3$-polytopes corresponding to nonselfcrossing Eulerian cycles on ideal right-angled hyperbolic $3$-polytopes.
Problem 4.
To characterise Hamiltonian cycles on simple $3$-polytopes corresponding to nonselfcrossing Eulerian cycles on $4$-valent convex polytopes.
Problem 5.
To characterise $4$-valent convex polytopes corresponding to Hamiltonian cycles on simple $3$-polytopes.
Problem 6.
To enumerate all combinatorially different nonselfcrossing Eulerian cycles on any ideal right-angled $3$-polytope.
6. Acknowledgements
The author is grateful to V.M. Buchstaber, A.A. Gaifullin, D.V. Gugnin, A.D. Mednykh, T.E. Panov, M.V. Prasolov, D.A. Tsygankov, A.Yu. Vesnin for fruitful discussions.