### v3

# On hyperbolic links associated to Eulerian subgraphs on right-angled hyperbolic $3$-polytopes of finite volume

Abstract.
We consider Eulerian cycles without transversal selfintersections in $4$-valent planar graphs. We prove that any cycle of this type in the graph of an ideal right-angled hyperbolic $3$-polytope corresponds to a hyperbolic link such that its complement consists of $4$-copies of this polytope glued according to its checkerboard coloring. Moreover, this link consists of trivially embedded circles bijectively corresponding to the vertices of the polytope. For such cycles we prove that the $3$-antiprism $A(3)$ (octahedron) has exactly $2$ combinatorially different cycles, the $4$-antiprism $A(4)$ has exactly $7$ combinatorially different cycles, and these cycles correspond to $7$ cycles (perhaps combinatorially equivalent) on any polytope different from antiprisms, and any antiprism $A(k)$ has at least $2$ combinatorially different cycles. The $2$-fold branched covering space corresponding to our link is a small cover over some simple $3$-polytope. This small cover is build from a Hamiltonian cycle on this polytope by the A.D. Mednykh’s construction. We show that any Hamiltonian cycle on a compact right-angled hyperbolic $3$-polytope arises in this way, while for a Hamiltonian cycle on a right-angled hyperbolic $3$-polytope of finite volume the necessary and sufficient condition is that at each ideal vertex it does not go straight. We introduce a transformation of a Eulerian cycle along conjugated vertices allowing to build new cycles from a given one. The link corresponding to a Hamiltonian cycle on a simple $3$-polytope always contains the Hopf link consisting of two circles. We consider links corresponding to Hamiltonian theta-graphs and Hamiltonian $K_{4}$-graphs on simple $3$-polytopes introduced by A.D. Mednykh and A.Yu. Vesnin. We give a criterion when such a link consists of mutually unlinked circles and when it is trivial. We give a necessary condition for such a link to be hyperbolic. The simplest example is the Borromean rings corresponding to the Hamiltonian theta-graph on the cube. We introduce the notions of a nonselfcrossing Eulerian theta-graph and $K_{4}$-graph on a right-angled hyperbolic $3$-polytope of finite volume with $2$ or $4$ finite vertices and construct the corresponding hyperbolic link.
Key words and phrases:
Convex polytope, Eulerian cycle, ideal right-angled hyperbolic polytope, small cover, Borromean rings, Brunnian link
2010 Mathematics Subject Classification:
57S12, 57S17, 57S25, 52B05, 57K10, 52B70, 57K10
1. Introduction
The theory of knot and links is a classical area of mathematics developing since XIX century. One of the well-known directions in this area is the theory of hyperbolic links. These are links with the complement having the structure of a complete hyperbolic manifold. In this paper using methods and results of toric topology we build a wide family of hyperbolic links corresponding to Eulerian cycles on any ideal right-angled hyperbolic $3$-polytope. Usually ideal right-angled polytopes arise when the alternating diagram of a link is reduced to some canonical form (see, for example [CKP21]). In our approach we build a link with trivial circles corresponding to vertices of the ideal right-angled polytope and their structure is defined by the Eulerian cycle. Any $k$-antiprism has a canonical Eulerian cycle. In this case our decomposition of the complement to the $(2k)$-link chain into $4$ antiprisms coincides with the decomposition described by W.P. Thurston [T02, Example 6.8.7].
Our main result (Theorem 3.2) is that any Eulerian cycle without transversal selfintersections in the graph of any ideal right-angled $3$-polytope $P$ corresponds to a link whose circles bijectively correspond to vertices of $P$ and the complement is decomposed into $4$ copies of $P$. We show that the $2$-fold branched covering space corresponding to this link is a small cover build by the A.D. Mednykh [M90] construction from a Hamiltonian cycle on a simple $3$-polytope. We show that any Hamiltonian cycle on a compact right-angled hyperbolic $3$-polytope arises in this way, while for a Hamiltonian cycle on a right-angled hyperbolic $3$-polytope of finite volume the necessary and sufficient condition is that at each ideal vertex it does not go straight. We provide methods how to build Eulerian cycles without transversal selfintersections on any ideal right-angled $3$-polytope.
The link corresponding to a Hamiltonian cycle on a simple $3$-polytope always contains the Hopf link consisting of two circles. In Section 7 we consider links corresponding to Hamiltonian theta-graphs and Hamiltonian $K_{4}$-graphs on simple $3$-polytopes introduced by A.D. Mednykh and A.Yu. Vesnin in [VM99S2]. We give a criterion when such a link consists of mutually unlinked circles and when it is trivial. We give a necessary condition for such a link to be hyperbolic. The simplest example is the Borromean rings corresponding to the Hamiltonian theta-graph on the cube.
In Section 8 we introduce the notions of a nonselfcrossing Eulerian theta-graph and $K_{4}$-graph on a right-angled hyperbolic $3$-polytope of finite volume with $2$ or $4$ finite vertices and construct the corresponding hyperbolic link.
2. Basic facts
Definition 2.1.
A cycle in a graph $G$ is called Eulerian if it passes each edge of the graph once (it may pass one vertex many times). We call an Eulerian cycle in the $4$-valent planar graph nonselfcrossing if at each vertex it does not intersect itself transversally, that is each time it visits the vertex it turns left or right, but does not go straight.
Remark 2.2.
In the paper [BFFS18] devoted to Barnette’s conjecture that every $3$-connected cubic planar bipartite graph is Hamiltonian, a nonselfcrossing Eulerian cycle is called an $A$-trail.
Remark 2.3.
As it was mentioned by D.V. Talalaev in private communication nonselfcrossing Eulerian cycles in $4$-valent planar graphs play an important role in the theory of electrical networks [BKT26].
A cycle in a graph is called Hamiltonian, if it visits each vertex once.
Let $G$ be a planar. Its medial graph is a new planar graph $M(G)$ with vertices bijectively corresponding to edges of $G$. Its edges arise when we walk around the boundary cycle of each face. Each vertex of this cycle corresponds to an edge of $M(G)$ connecting the vertices corresponding to successive edges of the cycle. A medial graph is $4$-valent. It is known that a graph $G$ is a graph $G(P)$ of an ideal hyperbolic right-angled $3$-polytope $P$ if and only if it is a medial graph of some simple polytope $\widehat{P}$. Moreover, $\widehat{P}$ is defined uniquely up to passing to the dual polytope $\widehat{P}^{*}$ (see more details in [E19]).
Proposition 2.4.
Nonselfcrossing Eulerian cycles in a $4$-valent planar graph $G$ correspond to Hamiltonian cycles in $M(G)$.
Construction 2.5 (A manifold from a checkerboard coloring).
Each ideal right-angled $3$-polytope $P$ admits a checkerboard coloring: its faces can be colored in black and white colors in such a way that adjacent faces have different colors (if $)G(P)=M(G(\widehat{P}$)), then black faces of $P$ correspond to vertices of $\widehat{P}$ and white facets of $P$ correspond to facets of $\widehat{P}$). Assign to white color the vector $e_{1}\in\mathbb{Z}_{2}^{2}$ and to black color $e_{2}\in\mathbb{Z}_{2}^{2}$. Then we obtain the mapping $\Lambda_{P}\colon\{F_{1},\dots,F_{m}\}\to\mathbb{Z}_{2}^{2}$, $F_{i}\to\Lambda_{i}$, from the set of facets of $P$ to $\mathbb{Z}_{2}^{2}$, and A.Yu.Vesnin–A.D. Mednykh construction (see [V17]) gives the complete hyperbolic manifold $N(P)$ of finite volume glued of $4$ copies of $P$:
$N(P)=P\times\mathbb{Z}_{2}^{2}/\sim,(p,a)\sim(q,b)\text{ if and only if }p=q\text{ and }a-b\in\langle\Lambda_{i}\colon p\in F_{i}\rangle.$
In this formula the ideal vertices are not assumed to belong to $P$.
It was proved in [E22b] that the family of manifolds $\{M(P)\}$, where $M(P)$ is the double of the manifold obtained from $N(P)$ by adding the boundary torus at each cusp, is cohomologically rigid over $\mathbb{Z}_{2}$, that is two manifolds from this family are homeomorphic if and only if their cohomology rings over $\mathbb{Z}_{2}$ are isomorphic as graded rings.
Definition 2.6.
A hyperelliptic manifold $M^{n}$ is an $n$-manifold with an action of an involution $\tau$ such that $M^{n}/\langle\tau\rangle$ is homeomorphic to $S^{n}$. The involution $\tau$ is called hyperelliptic.
In [M90] A.D. Mednykh constructed examples of hyperelliptic $3$-manifolds with geometric structures modelled on five of eight Thurston’s geometries: $\mathbb{R}^{3}$, $\mathbb{L}^{3}$, $\mathbb{S}^{3}$, $\mathbb{L}^{2}\times\mathbb{R}$, and $\mathbb{S}^{2}\times\mathbb{R}$. Each example was built using a right-angled $3$-polytope $P$ equipped with a Hamiltonian cycle. This construction can be described as follows.
Construction 2.7 (A small cover and a link from a Hamiltonian cycle).
Let $\Gamma$ be a Hamiltonian cycle in the graph of a simple $3$-polytope $Q$. Then it divides $\partial Q\simeq S^{2}$ into two disks. Each edge of $Q$ not lying in $\Gamma$ divides one of the disks into two disks. Thus, the adjacency graph of faces of $Q$ lying in the closure of each component of $\partial Q\setminus\Gamma$ is a tree and these faces can be colored in two colors in such a way that adjacent faces have different colors. Then combining both components we obtain a coloring of faces of $Q$ in four colors. This coloring corresponds to a mapping $\widetilde{\Lambda}_{\Gamma}$ of the set of faces of $Q$ to $\mathbb{Z}_{2}^{3}$ by the rule: the first three colors correspond to basic vectors $e_{1},e_{2},e_{3}\in\mathbb{Z}_{2}^{3}$, and the fourth corresponds to their sum $e_{1}+e_{2}+e_{3}$. This mapping gives an orientable $3$-manifold $N(Q,\widetilde{\Lambda}_{\Gamma})$ (it toric topology [BP15] it is called an orientable small cover):
$N(Q,\widetilde{\Lambda}_{\Gamma})=Q\times\mathbb{Z}_{2}^{3}/\sim,(p,a)\sim(q,b)\text{ if and only if }p=q\text{ and }a-b\in\langle\Lambda_{i}\colon p\in F_{i}\rangle.$
If the components of $\partial Q\setminus\Gamma$ correspond to the pairs of colors $(1,2)$ and $(3,4)$, then $\tau=e_{1}+e_{2}=e_{3}+(e_{1}+e_{2}+e_{3})$ is the hyperelliptic involution on $N(Q,\widetilde{\Lambda}_{\Gamma})$. Moreover, the mapping $N(P,\widetilde{\Lambda}_{\Gamma})\to S^{3}$ is a $2$-fold branched covering with the following branch set (see details in [EE25, Section 4.5]). The edges of $Q$ not lying in $\Gamma$ form a matching $M_{\Gamma}$ of the graph $G(Q)$ – a disjoint set of edges. For Hamiltonian cycle this matching is perfect, that is it covers all the vertices (in graph theory perfect matchings are also called $1$-factors). Then the preimage of $M_{\Gamma}$ in $N(Q,\widetilde{\Lambda}_{\Gamma})$ and in $S^{3}$ is a disjoint set of circles $C_{\Gamma}$, each circle glued of two copies of the corresponding edge. This link is the branch set of the covering. The detailed description of this link is given in [G24]. In [VM99S2] this construction was generalized from Hamiltonian cycles to Hamiltonian theta-graphs and Hamiltonian $K_{4}$-subgraphs. We will discuss the corresponding links in Section 7.
On the language of toric topology this construction and its generalizations is described in [E24] and [EE25]. In particular, in these papers the space corresponding to general vector-coloring of rank $r$ is considered.
Definition 2.8.
A vector-coloring of rank $r$ of a simple $3$-polytope $Q$ is a mapping $\Lambda$ from the set of its facets $F_{1},\dots,F_{m}$ to $\mathbb{Z}_{2}^{r}$, $F_{i}\to\Lambda_{i}$, such that $\langle\Lambda_{1},\dots,\Lambda_{m}\rangle=\mathbb{Z}_{2}^{r}$. It corresponds to the space
$N(Q,\Lambda)=Q\times\mathbb{Z}_{2}^{r}/\sim,(p,a)\sim(q,b)\text{ if and only if }p=q\text{ and }a-b\in\langle\Lambda_{i}\colon p\in F_{i}\rangle.$
This space has an action of $\mathbb{Z}_{2}^{r}$. This space is a manifold if and only if for each vertex $v=F_{i}\cap F_{j}\cap F_{k}$ either $\Lambda_{i}=\Lambda_{j}=\Lambda_{k}$, or two of these vectors are equal and the third is different, or three vectors are linearly independent (see [E24]). In particular, any simple cycle $\Gamma$ in the graph of $Q$ divides $\partial Q$ in two connected components. Define $\Lambda_{i}=e_{1}$, if $F_{i}$ lies in the closure of the first component, and $\Lambda_{i}=e_{2}$, if $F_{i}$ lies in the closure of the second component. This vector-coloring $\Lambda_{\Gamma}$ has rank $2$ and $N(Q,\Lambda_{\Gamma})\simeq S^{3}$ glued of $4$ copies of $Q$ in the following way. $\partial Q$ is divided by $\Gamma$ into two facets. The complex on the polytope $Q$ with these two facets is homeomorphic to the $3$-ball with the boundary divided into two hemispheres by the equator. Then the ball $Q\times(0,0)$ is glued to $Q\times(1,0)$ along the hemisphere corresponding to $e_{1}$ to give a new ball, as well as $Q\times(0,1)$ to $Q\times(1,1)$, while the resulting balls are glued along the boundaries to give the $3$-sphere.
3. Main construction
Construction 3.1.
Let $\gamma$ be a nonselfcrossing Eulerian cycle on the $4$-valent 33-polytope $P$. Let us build a simple $3$-polytope $Q(P,\gamma)$ with a Hamiltonian cycle $\Gamma_{\gamma}$ by the following rule. Substitute each vertex of the graph of $P$ by two vertices connected by an edge in such a way that each pair of successive edges of $\gamma$ at this vertex is incident to the same vertex of the new edge. The new graph satisfies the condition that each face is bounded by a simple edge-cycle, and if two boundary cycles of faces intersect, then by an edge. Thus by the Steinitz theorem this graph is a graph of a unique combinatorial simple polytope $Q(P,\gamma)$ (see more details in [E19]). Moreover, $\gamma$ corresponds to a Hamiltonian cycle $\Gamma_{\gamma}$ on this polytope. The new edges form a perfect matching in the graph of $Q$.
Theorem 3.2.
Let $P$ be an ideal right-angled hyperbolic $3$-polytope. Each nonselfcrossing Eulerian cycle $\gamma$ corresponds to a link $C_{\gamma}\subset S^{3}$ consisting of $\{\text{vertices of }P\}$ circles such that its complement $S^{3}\setminus C_{\gamma}$ is homeomorphic to the hyperbolic manifold $N(P)$ glued of $4$ copies of $P$. Moreover, $C_{\gamma}$ is the branch set of the $2$-fold branched covering $N(Q(P,\gamma),\widetilde{\Lambda}_{\Gamma_{\gamma}})\to S^{3}$.
Proof.
Indeed, the $2$-fold branched covering has the form $N(Q,\widetilde{\Lambda}_{\Gamma_{\gamma}})\to N(Q,\Lambda_{\Gamma_{Q}})$: $p,a]\to[p,\pi(a)]$, where $\pi\colon\mathbb{Z}_{2}^{3}\to\mathbb{Z}_{2}^{3}/\langle e_{1}+e_{2}\rangle\simeq\mathbb{Z}_{2}^{2}$ is a projection. Under the homeomorphism $N(Q,\Lambda_{\Gamma_{Q}})\simeq S^{3}$ the preimage of the perfect matching on $Q$ corresponds to the link $C_{\gamma}$ consisting of $\{\text{vertices of }P\}$ circles (each circle corresponds to an edge of the perfect matching and to a vertex of $P$). $Q\setminus\{\text{perfect matching}\}$ is homeomorphic to $P$, while $S^{3}\setminus C_{\gamma}$ is homeomorphic to $N(P)$. ∎
Remark 3.3.
A link $L$ is called hyperbolic if its complement $S^{3}\setminus L$ has a structure of a complete hyperbolic manifold. In [CKP21] a hyperbolic link was called right-angled, if $S^{3}\setminus L$ with the complete hyperbolic structure admits a decomposition into ideal hyperbolic right-angled polytopes. By construction the link $C_{\gamma}$ is right-angled.
Remark 3.4.
For nonselfcrossing Eulerian cycles on any $4$-valent convex $3$-polytope the analog of Theorem 3.2 without hyperbolic structure on $S^{3}\setminus C_{\gamma}$ holds.
Remark 3.5.
The transition from a Hamiltonian cycle $\Gamma$ on a cubic (that is $3$-valent) graph $G$ to the nonselfcrossing Eulerian cycle in the graph $G/M_{\Gamma}$ obtained by contracting the edges of the perfect matching $M_{\Gamma}$ was used in [BP17] to give a new characterisation of cubic Hamiltonian graphs having a perfect matching.
Example 3.6.
The octahedron is a unique right-angled polytope with the smallest number of vertices (equal to $6$). Up to combinatorial symmetries it has exactly two nonselfcrossing Eulerian cycles (see the proof in Fig. 1) shown in Fig. 2. We also present the corresponding simple polytopes and hyperbolic links. These links are not isotopic, but their complements are homeomorphic. The ways to represent links follows the method from [T25].
Figure 1. Enumeration of nonselfcrossing Eulerian cycles on the octahedron
Figure 2. Hyperbolic links corresponding to nonselfcrossing Eulerian cycles on the octahedron
Example 3.7.
Example 3.6 can be generalized as follows. It is known that any antiprism $A(k)$ ($A(3)$ is the octahedron) is an ideal right-angled $3$-polytope (see [V17]). In Fig 3 we show two different nonselfcrossing Eulerian cycles on this polytope. The hyperbolic structure on the complement to the link corresponding to the left cycle is exactly the structure defined by W.P.Thurston in the complement to the $(2k)$-link chain [T02, Example 6.8.7]. We learn this example due to the lectures by A.Yu.Vesnin and the diploma works by D.V. Chepakova [C23] and D.A. Tsygankov [T25]. The case of $A(3)$ is also mentioned in [V17, Section 5.1].
Figure 3. Nonselfcrossing Eulerian cycles on the antiprism
Example 3.8.
It can be shown (see Fig. 4) that up to combinatorial symmetries $A(4)$ has exactly 77 nonselfcrossing Eulerian cycles shown in Fig. 5.
Figure 4. Enumeration of nonselfcrossing Eulerian cycles on $A(4)$
Figure 5. Nonselfcrossing Eulerian cycles on $A(4)$ and the corresponding polytopes
4. Edge-twists and the existence of nonselfcrossing Eulerian cycles
In [V17, Theorem 2.14] on the base of results from [BGGMTW05] the following construction of all the ideal right-angled polytopes was described .
Definition 4.1.
An operation of an edge-twist is shown in Fig. 6. Two edges on the left belong to one facet of a the ideal right-angled polytope and connect $4$ distinct vertices. The result is again an ideal right-angled polytope. Let us call an edge-twist restricted if both edges are adjacent to the same edge, that is the $4$ vertices follow each other during the round walk along the boundary of a facet.
Figure 6. An operation of an edge-twist
Theorem 4.2 ([V17]).
Any ideal right-angled $3$-polytope can be obtained by operations of an edge-twist from some $k$-antiprism $A(k)$, $k\geqslant 3$.
Remark 4.3.
Operations of an edge-twist are not applicable to the octahedron, hence all the other polytopes are obtained from $k$-antiprisms, $k\geqslant 4$.
In [E19, Theorem 9.13] this result was improved.
Theorem 4.4 ([E19]).
A $3$-polytope is an ideal right-angled $3$-polytope if and only if either it is a $k$-antiprism $A(k)$, $k\geqslant 3$, or it can be obtained from the $4$-antiprism by operations of a restricted edge-twist.
The following result is straightforward from the definitions.
Proposition 4.5.
Any edge-twist transforms a nonselfcrossing Eulerian cycle to a nonselfcrossing Eulerian cycle on the new polytope.
Corollary 4.6.
The $3$-antiprism $A(3)$ (octahedron) has exactly $2$ combinatorially different nonselfcrossing Eulerian cycles, the $4$-antiprism $A(4)$ has exactly $7$ combinatorially different nonselfcrossing Eulerian cycles, and they correspond to $7$ nonselfcrossing Eulerian cycles (perhaps some of them are combinatorially equivalent) on any polytope different from antiprisms, and any antiprism $A(k)$ has at least $2$ combinatorially different cycles.
Proof.
This follows from Examples 3.6, 3.7, and 3.8. ∎
Remark 4.7.
As was mentioned by A.A. Gaifullin, the existence of a nonselfcrossing Eulerian cycle can be proved as follows. Since each vertex of $P$ has even valency, it has a Eulerian cycle. Then we can deform this cycle at bad vertices. If it has a transversal self-crossing, then one of the two ways to change this crossing leaves the cycle connected.
Problem 1.
To enumerate all combinatorially different nonselfcrossing Eulerian cycles on any ideal right-angled $3$-polytope. To find estimates for their number.
5. Transformations of nonselfcrossing Eulerian cycles
Definition 5.1.
We will call two edges $E_{1}$ and $E_{2}$ of $Q$ not lying in a Hamiltonian cycle $\Gamma$ conjugated, if each edge intersects both components of the complement in $\Gamma$ to the vertices of the other edge (in other words, if $\Gamma\cup E_{1}\cup E_{2}$ is homeomorphic to the full graph $K_{4}$ on four vertices). We call two vertices of an ideal right-angled $3$-polytope $P$ conjugated along the nonselfcrossing Eulerian cycle $\gamma$, if the corresponding edges of $Q(P,\gamma)$ are conjugated.
Proposition 5.2.
The circles in $C_{\Gamma}$ corresponding to the edges of $Q$ not lying in $\Gamma$ are linked if and only if the edges are conjugated.
Proof.
This becomes evident if we look at the link in a way shown in Fig. 2 on the right. ∎
Lemma 5.3.
Each edge of $Q\setminus\Gamma$ has a conjugated edge.
Proof.
Indeed, let the edge $E$ of $Q\setminus\Gamma$ have no conjugated edges. $E$ is the intersection of two facets $F_{i}$ and $F_{j}$ of $Q$ lying in the closure of the same connected component of $\partial Q\setminus\Gamma$. Then both vertices of $E$ belong to the same facet $F_{k}$ lying in the closure of the other connected component. In this case $E$ belongs to $F_{k}$, which is a contradiction. ∎
Corollary 5.4.
Each circle of the link $C_{\Gamma}$ corresponding to a Hamiltonian cycle $\Gamma$ on a simple $3$-polytope $Q$ is linked to at least one other circle of $C_{\Gamma}$.
Construction 5.5 (Transformation of a nonselfcrossing Eulerian cycle along conjugated vertices).
Given two conjugated vertices $v$ and $w$ of a nonselfcrossing Eulerian cycle $\gamma$ of an ideal right-angled $3$-polytope $P$ we can build a new nonselfcrossing Eulerian cycle in the following way. In both vertices we change the pairs of successive edges of the cycle to complementary pair (see Fig. 7). In Fig. 8 we show how the Hamiltonian cycle $\Gamma_{\gamma}$ is transformed under this operation (the new Hamiltonian cycle belongs to another polytope obtained from $Q$ by two flips).
Figure 7. Local transformation of the Eulerian cycle $\gamma$ and the corresponding flip of the polytope $Q$
Figure 8. Transformation of the Hamiltonian cycle $\Gamma_{\gamma}$.
Proposition 5.6.
Let $\gamma$ be a nonselfcrossing Eulerian cycle on the ideal right-angled $3$-polytope $P$. Then for any vertex of $P$ there is at least one conjugated vertex and the corresponding transformation of $\gamma$.
Conjecture 5.7.
Any two nonselfcrossing Eulerian cycles are connected by a sequence of transformations along conjugated vertices.
6. Links associated to Hamiltonian cycles on right-angled $3$-polytopes
A $k$-belt is a cyclic sequence of $k$ facets such that facets are adjacent if and only if they are successive and no three facets have a common vertex. It follows from results by A.V. Pogorelov and E.M. Andreev that a $3$-polytope is combinatorially equivalent to a compact right-angled hyperbolic $3$-polytope if and only if it is a simple polytope different from the simplex and has no $3$- and $4$-belts (see more details in [E19]). These polytopes are called Pogorelov polytopes.
Proposition 6.1.
Every compact right-angled hyperbolic $3$-polytope $Q$ with a Hamiltonian cycle $\Gamma$ defines an ideal right-angled polytope $P$ obtained by shrinking to points the edges of the perfect matching consisting of the edges not in the cycle. The polytope $Q$ has an induced nonselfcrossing Eulerian cycle $\gamma$ such that $\Gamma=\Gamma_{\gamma}$. For the link $C_{\gamma}$ both the complement $S^{3}\setminus C_{\gamma}$ and the $2$-fold branched covering space have complete hyperbolic structures obtained by gluing right-angled polytopes.
Proof.
Indeed, by [E19, Theorem 11.6] the polytope obtained by cutting off all these edges is almost Pogorelov, and by [E19, Theorem 6.5] shrinking the obtained quadrangles to points give the ideal right-angled $3$-polytope. ∎
Remark 6.2.
In [BD25] the contraction of edges of right-angled hyperbolic polytopes producing right-angled polytopes of finite volume is discussed. In particular, the antiprism $A_{n}$ is obtained by a contraction of a perfect matching of the Löbell polytope ($n$-barrel) $L_{n}$. The inverse operation corresponds to the hyperbolic Dehn filling.
Example 6.3.
The dodecahedron is a unique compact right-angled polytope with minimal number of facets (equal to $12$). Up to combinatorial symmetries it has a unique Hamiltonian cycle. In Fig. 9 we show this Hamiltonian cycle, and the way how the associated ideal right-angled polytope is obtained from $A(4)$ by a sequence of two restricted edge-twists. We also show another polytope corresponding to another Eulerian cycle on the $4$-antiprism. The corresponding links have homeomorphic complements, but the first link has the $2$-fold branched covering space with a hyperbolic structure, and the $2$-fold branched covering space of the second link contains incompressible tori corresponding to $4$-belts (see more details in [E22a]).
Figure 9. Hamiltonian cycle on the dodecahedron
Example 6.4.
A fullerene is a simple $3$-polytope with only pentagonal and hexagonal faces. It is known that any fullerene is a right-angled hyperbolic polytope and the dodecahedron is the fullerene with minimal number of facets (see more details in [E19]). It was shown by F. Kardoš [K20] that any fullerene has a Hamiltonian cycle. Each Hamiltonian cycle on a fullerene corresponds to a hyperbolic link with hyperbolic $2$-fold branched covering space.
Remark 6.5.
For Hamiltonian cycles on general simple $3$-polytopes the analog of Proposition 6.1 is not valid. If we shrink all the edges of the perfect matching complementary to the cycle to points, the resulting graph may not represent a polytope. For example, if $Q$ contains a triangle, then the resulting spherical complex contains a bigon. For a Hamiltonian cycle on the cube the resulting complex also contains a bigon. But if the resulting spherical graph contains no bigonal faces (for example when $Q$ has no triangles and quadrangles), then by the Steinitz theorem it is the graph of a $4$-valent convex polytope. But this polytope may be not ideal right-angled.
Now consider right-angled hyperbolic $3$-polytopes of finite volume. Each finite vertex of such a polytope has valency $3$, and ideal vertices have valency $4$. Using Andreev’s theorem it can be shown that cutting off ideal vertices defines a bijection between combinatorial types of right-angled hyperbolic $3$-polytopes of finite volume and almost Pogorelov polytopes different from the $4$-prism (cube) and the $5$-prism (see [DO01, Theorem 10.3.1] and [E19, Theorem 6.5]). Moreover, all quadrangles of the resulting polytope arise from ideal vertices. A simple $3$-polytope is called an almost Pogorelov polytope, if it is different from the simplex, has no $3$-belts, and any its $4$-belt surrounds a quadrangular facet. The simple polytope with $8$ facets drawn in the center at the bottom in Fig. 2 we will denote $P_{8}$. It has a nontrivial $4$-belt consisting of pentagons and surrounding two quadrangles on each side. It was shown in [E19] that this polytope has some properties similar to properties of almost Pogorelov polytopes.
Proposition 6.6.
Let $\Gamma$ be a Hamiltonian cycle on an almost Pogorelov $3$-polytope $Q$ or the polytope $P_{8}$.
  1. (1)
    Then shrinking to points the edges of the perfect matching $M_{\Gamma}$ complementary to $\Gamma$ gives an ideal right-angled polytope $P$ if and only if each quadrangle of $Q$ has three edges in $\Gamma$. In this case the polytope $Q$ has an induced nonselfcrossing Eulerian cycle $\gamma$ such that $\Gamma=\Gamma_{\gamma}$. For the link $C_{\gamma}$ the complement $S^{3}\setminus C_{\gamma}$ is hyperbolic.
  2. (2)
    The cube $I^{3}$ and the $5$-prism $M_{5}\times I$ do not have Hamiltonian cycles with the above condition. The polytope $P_{8}$ up to combinatorial symmetries has a unique Hamiltonian cycle with the above condition shown in Fig. 2.
  3. (3)
    For $Q\notin\{I^{3},M_{5}\times I,P_{8}\}$ the $2$-fold branched covering space corresponding to $C_{\Gamma}$ becomes hyperbolic after cutting along incompressible Klein bottles corresponding to quadrangles of $Q$. For $Q=Q_{8}$ the $2$-fold branched covering space splits into two manifolds with geometry $\mathbb{L}^{2}\times\mathbb{R}$ after cutting along the incompressible torus corresponding to the $4$-belt consisting of pentagons.
Proof.
For each quadrangle of $Q$ the Hamiltonian cycle contains either two its opposite edges, or three of its edges. In the first case after shrinking the quadrangle becomes a bigon and we obtain no polytope. If each quadrangle intersects $\Gamma$ by three edges, then it intersects at vertices two edges of the complementary perfect matching, and by  [E19, Corollary 12.31] cutting off all the edges of the matching gives an almost Pogorelov polytope different from the $4$- and the $5$-prism and produces all its quadrangles. For $Q=P_{8}$ to use [E19, Corollary 12.31] we additionally need to check that $\Gamma$ does not contain all the four common edges of pentagons. But in this case it can not have $3$ edges on each quadrangle, so the additional condition holds. Now shrinking quadrangles to points we obtain an ideal right-angled hyperbolic $3$-polytope $P$. The same polytope we obtain by shrinking all the edges of the matching. Thus, item (1) is proved.
Item (2) follows from the direct enumeration of Hamiltonian cycles on $I^{3},M_{5}\times I$ and $P_{8}$.
By [E22a, Theorem 4.12] for $Q\notin\{I^{3},M_{5}\times I,P_{8}\}$ its quadrangles correspond to incompressible Klein bottles in $N(Q,\widetilde{\Lambda}_{\Gamma})$ such that the complement to their union has a complete hyperbolic structure of finite volume. Also by this Theorem for $Q=P_{8}$ the $4$-belt consisting of pentagons corresponds to an incompressible torus in $N(Q,\widetilde{\Lambda}_{\Gamma})$ such that its complement consists of two manifolds with geometry $\mathbb{L}^{2}\times\mathbb{R}$. This proves item (3). ∎
Corollary 6.7.
A Hamiltonian cycle on a right-angled hyperbolic $3$-polytope of finite volume corresponds to a nonselfcrossing Eulerian cycle on the ideal right-angled hyperbolic $3$-polytope if and only if at each ideal vertex it turns left or right, but does not go straight.
Example 6.8.
In Fig. 10 we show a Hamiltonian cycle on the $3$-dimensional permutohedron intersecting each quadrangle by $3$ edges. After shrinking quadrangles to points we obtain a Hamiltonian cycle on the ideal octahedron.
Figure 10. The Hamiltonian cycle on the permutohedron corresponding to the Hamiltonian cycle on the ideal octahedron.
Problem 2.
To characterise ideal right-angled $3$-polytopes corresponding to Hamiltonian cycles on (a) compact right-angled hyperbolic $3$-polytopes (b) right-angled hyperbolic $3$-polytopes of finite volume.
Remark 6.9.
In [E19, Theorem 9.17] is was proved that any ideal right-angled hyperbolic $3$-polytope $P$ can be obtained from an almost Pogorelov polytope or the polytope $P_{8}$ by a contraction of edges of a perfect matching such that no quadrangle contains two edges of the matching. Nevertheless, the complement to this matching may be not a Hamiltonian cycle, but a union of cycles containing all the vertices of the polytope.
Problem 3.
To characterise Hamiltonian cycles on simple $3$-polytopes corresponding to nonselfcrossing Eulerian cycles on (a) ideal right-angled hyperbolic $3$-polytopes (b) $4$-valent convex polytopes.
Problem 4.
To characterise $4$-valent convex polytopes corresponding to Hamiltonian cycles on simple $3$-polytopes.
7. Links corresponding to Hamiltonian theta-graphs and $K_{4}$-graphs
This section arose due to the question posed by Victor Buchstaber: using technique of toric topology to build a rich family of Brunnian links that is nontrivial links that become a set of trivial unlinked circles if any one component is removed. The question is motivated by the notion of a Efimov state [E70] in quantum mechanics. This is a bound state of three bosons such that the two-particle attraction is too weak to allow two bosons to form a pair. If one of the particles is removed, the remaining two fall apart. The Efimov sate is symbolically depicted by the Borromean rings, which is the first nontrivial example of a Brunnian link. This link arises in our considerations and corresponds to a Hamiltonian theta-graph on the cube. It turned out that using methods from [VM99S2] we can naturally build a family of links that are not Brunnian, but each link consists of trivial pairwise unlinked circles. Moreover, if the link is nontrivial, then it contains the Borromean rings.
Remark 7.1.
In algebraic topology the Borromean rings are associated to the triple Massey product – an operation producing a new cohomology class from three classes with trivial pairwise products. In particular, for the complement of the Borromean rings the triple Massey product is defined and non-zero [M68]. The product of the $1$-cochains dual to the $3$ rings via Alexander duality is zero, while the triple Massey product is non-zero. As we will see in Example 7.7, toric topology associates to the Borromean rings also a compact $12$-dimensional manifold with a nontrivial triple Massey product. It is the moment-angle manifold of the $3$-dimensional associahedron.
Remark 7.2.
In [RV25] a family of hyperbolic Brunnian links was constructed starting from links $L_{3n+2}$ consisting of $3n+2$ components with the complement $S^{3}\setminus L_{3n+2}$ decomposed into $4$ right-angled hyperbolic $(2n)$-antiprisms $A_{2n}$.
Construction 7.3 (Hyperelliptic manifold and link from a Hamiltonian theta- or $K_{4}$-graph).
In the language of toric topology the Mednykh-Vesnin construction from [VM99S2] can be described as follows (see also [E24] and [EE25]). A Hamiltonian theta-graph in the simple $3$-polytope $Q$ is a subset of the graph $G(Q)$ consisting of three disjoint simple paths connecting two vertices of $Q$ and containing all the vertices of $Q$. A Hamiltonian $K_{4}$-graph in $Q$ is a subset of $G(Q)$ consisting of $4$ vertices and $6$ disjoint paths connecting any two of these vertices and containing all the vertices of $Q$.
Let $\Gamma$ be a Hamiltonian theta- or $K_{4}$-graph of a simple $3$-polytope $Q$. Then it divides $\partial Q\simeq S^{2}$ into $k=3$ (for theta-graph) or $k=4$ (for $K_{4}$-graph) disks. Each edge of $Q$ not lying in $\Gamma$ divides one of the disks into two disks. Thus, the adjacency graph of faces of $Q$ lying in the closure of each component of $\partial Q\setminus\Gamma$ is a tree and these faces can be colored in two colors (black and white) in such a way that adjacent faces have different colors. Let $a_{1}$, $\dots$, $a_{k}$, $b_{1}$ be a basis in $\mathbb{Z}_{2}^{k+1}$. Define $b_{2}$, $\dots$, $b_{k}$ by the rule $a_{i}+b_{i}=\tau=a_{1}+b_{1}$ for all $i$. Assign to each facet of $Q$ in $i$-th component of $\partial Q\setminus\Gamma$ the vector $a_{i}$ if it is white and $b_{i}$ if it is black. We obtain the vector-coloring $\widetilde{\Lambda}_{\Gamma}$ of rank $k+1$ and the orientable manifold $N(Q,\widetilde{\Lambda}_{\Gamma})$ with the action of $\mathbb{Z}_{2}^{k+1}$. Then $\tau$ is a hyperelliptic involution and $N(Q,\widetilde{\Lambda}_{\Gamma})/\langle\tau\rangle=N(Q,\Lambda_{\Gamma})\simeq S^{3}$, where $\Lambda_{\Gamma}$ is the composition $\pi\circ\widetilde{\Lambda}_{\Gamma}$, where $\pi\colon\mathbb{Z}_{2}^{k+1}\to\mathbb{Z}_{2}^{k+1}/\langle\tau\rangle\simeq\mathbb{Z}_{2}^{k}$.
The homeomorphism $N(Q,\Lambda_{\Gamma})\simeq S^{3}$ can be seen as follows. All facets of $i$-th connected component of $Q\setminus\Gamma$ are colored in the same vector $a_{i}]\in\mathbb{Z}_{2}^{k}$, and the vectors $a_{1}],\dots,[a_{k}]$ form a basis in $\mathbb{Z}_{2}^{k}$. For the theta-graph $\Gamma$ let $v$ and $w$ be its vertices. Then there is a homeomorphism of $Q\setminus w$ to the positive octant in $\mathbb{R}^{3}$ mapping $v$ to the origin and the paths to coordinate rays. Then the space $N(Q,\Lambda_{\Gamma})\setminus[w\times\mathbb{Z}_{2}^{k}/\sim]$ is equivariantly homeomorphic to $\mathbb{R}^{3}$ with the involution $a_{i}]$ corresponding to the change of sign of the $i$-th coordinate.
For the $K_{4}$-graph $\Gamma$ the complex on $Q$ given by edges and faces of this graph is homeomorphic to the face complex of the simplex $\Delta^{3}$, and $N(Q,\Lambda_{\Gamma})$ is equivariantly homeomorphic to the real moment-angle manifold $\mathbb{R}\mathcal{Z}_{\Delta^{3}}\simeq S^{3}$. It can be visualised similarly as for the theta-graph. Namely, there is a homeomorphism of $Q$ to the simplex $\Delta^{3}$ that is the convex hull of the origin and the ends of the three basis vectors. Then the vectors corresponding to three coordinate facets correspond to reflections in these facets. Gluing $8$ copies of $\Delta^{3}$ we obtain the octahedron $Oct^{3}$. Also for each octant the complement to the reflected copy of $\Delta^{3}$ is homeomorphic to $\Delta^{3}\setminus\{\text{Origin}\}\simeq Q\setminus\{v\}$, where $v$ is a vertex of $\Gamma$. Then these complements are glued to $\mathbb{R}^{3}\setminus Oct^{3}$.
The mapping $N(Q,\widetilde{\Lambda}_{\Gamma})\to N(Q,\Lambda_{\Gamma})\simeq S^{3}$ is a $2$-fold branched covering with the following branch set (see details in [EE25, Section 4.5]). The edges of $Q$ not lying in $\Gamma$ form a matching $M_{\Gamma}$ of $G(Q)$ and the preimage of this set in $N(Q,\widetilde{\Lambda}_{\Gamma})$ and in $S^{3}$ is a disjoint set of circles $C_{\Gamma}$. This link is the branch set of the covering. For $k=3$ each edge of $M_{\Gamma}$ corresponds either to a circle glued of $4$ copies of this edge (if the edge has vertices on different paths of $\Gamma$), or to a pair of circles each glued of $2$ copies of the edge (if the edge has vertices on the same path of $\Gamma$). For $k=4$ it corresponds either to a pair of circles glued of $4$ copies of the edge (if the edge has vertices on different paths of $\Gamma$), or to $4$ circles each glued of $2$ copies of the edge (if the edge has vertices on the same path of $\Gamma$).
Example 7.4.
In Fig.  11 we show the link corresponding to a Hamiltonian theta-graph on the cube. It is the Borromean rings. The manifold $N(Q,\widetilde{\Lambda}_{\Gamma})$ has a Euclidean structure.
Example 7.5.
In Fig. 12 we show the link corresponding to a Hamiltonian theta-graph on the dodecahedron. The manifold $N(Q,\widetilde{\Lambda}_{\Gamma})$ has a hyperbolic structure.
Figure 11. The Borromean rings corresponding to a Hamiltonian theta-graph on the cube
Figure 12. The link corresponding to a Hamiltonian theta-graph on the dodecahedron
We have the following generalization of Proposition 6.6.
Proposition 7.6.
Let $\Gamma$ be a Hamiltonian theta-graph or $K_{4}$-graph on an almost Pogorelov $3$-polytope $Q$ or the polytope $P_{8}$. Then shrinking to points the edges of the matching $M_{\Gamma}$ complementary to $\Gamma$ gives a right-angled hyperbolic polytope $P$ of finite volume (with $2$ proper vertices for theta-graph and $4$ proper vertices for $K_{4}$-graph) if and only if each quadrangle of $Q$ intersects at a vertex at least one edge of $M_{\Gamma}$, and if $Q=P_{8}$ then additionally at least one common edge of pentagons belongs to $M_{\Gamma}$. For the link $C_{\Gamma}$ the complement $S^{3}\setminus C_{\Gamma}$ is hyperbolic.
Proof.
By [E19, Corollary 12.31] cutting off all the edges of the matching gives an almost Pogorelov polytope different from the $4$- and the $5$-prism and produces all its quadrangles. Then shrinking quadrangles to points we obtain a right-angled polytope $P$ of finite volume. The same polytope we obtain by shrinking all the edges of the matching. ∎
Example 7.7.
The theta-graph from Example 7.4 satisfies conditions of Proposition 7.6. The polytope $P$ is a right-angled $3$-gonal bipyramid with $2$ proper and $3$ ideal vertices. The complement of the Borromean rings $C_{\Gamma}$ is glued of $8$ copies of $P$. The almost Pogorelov polytope associated to the $3$-gonal bipyramid is the $3$-dimensional associahedron (Stasheff polytope) $As^{3}$. As is it mentioned in [BP15, Remark after Example 4.9.4] the $12$-dimensional moment-angle manifold $\mathcal{Z}_{As^{3}}$ has a nontrivial triple Massey product (it follows also from [DS07, Theorem 6.1.1]). In [L19] nontrivial triple Massey products were constructed in cohomology of moment-angle manifolds of more general graph-associahedra.
Corollary 7.8.
Any Hamiltonian theta-graph or $K_{4}$-graph $\Gamma$ on the compact right-angled hyperbolic $3$-polytope $Q$ defines a right-angled hyperbolic polytope $P$ of finite volume (with $2$ or $4$ proper vertices) obtained by shrinking to points the edges of $M_{\Gamma}$. For the link $C_{\Gamma}$ both the complement $S^{3}\setminus C_{\Gamma}$ and the $2$-fold branched covering space have complete hyperbolic structures obtained by gluing right-angled polytopes.
Example 7.9.
The theta-graph from Example 7.5 gives the right-angled polytope $P$ with $2$ proper and $9$ ideal vertices. The complement of $C_{\Gamma}$ is glued of $8$ copies of $P$.
It is easy to see that each link $C_{\Gamma}$ corresponding to a Hamiltonian cycle, theta-graph or $K_{4}$-graph on a simple $3$-polytope $Q$ consists of trivial circles. Moreover, as it was shown in Corollary 5.4 in the case of a Hamiltonian cycle on a simple $3$-polytope each circle is linked to at least one other circle, in particular $C_{\Gamma}$ is nontrivial.
Construction 7.10 (Cutting off a vertex of $\Gamma$).
[…]
Figure 13. Cutting off a vertex
Example 7.11.
Let $\Gamma_{0}$ be the Hamiltonian theta-graph on the simplex $\Delta^{3}$, obtained by deletion of any edge from the graph $G(\Delta^{3})=K_{4}$. Up to combinatorial symmetries it is a unique theta-graph in $G(\Delta^{3})$. The link $C_{\Gamma_{0}}$ is a trivial circle. Then for any pair $(Q,\Gamma)$ obtained from $(\Delta^{3},\Gamma_{0})$ by a sequence of operations of cutting off a vertex the corresponding link $C_{\Gamma}$ is trivial.
Theorem 7.12.
Let $\Gamma$ be a Hamiltonian theta-graph on a simple $3$-polytope $Q$. Then
  1. (1)
    the link $C_{\Gamma}$ consists of mutually unlinked circles if and only if each edge of $M_{\Gamma}$ connects vertices of different paths of $\Gamma$;
  2. (2)
    if $C_{\Gamma}$ consists of mutually unlinked circles and is nontrivial, then it contains a triple of Borromean rings;
  3. (3)
    the link $C_{\Gamma}$ is trivial if and only if $(Q,\Gamma)$ is obtained from $(\Delta^{3},\Gamma_{0})$, by a sequence of operations of cutting off a vertex.
Example 7.13.
The link in Example 7.5 consists of mutually unlinked circles and contains many triples of Borromean rings.
Proof of Theorem 7.12.
From the representation of the theta-graph on the coordinate rays of the octant it is clear that each two circles are unlinked if each edge of the matching connects two different paths of $\Gamma$. On the other hand, if there is an edge $M_{\Gamma}$ connecting two vertices $v$ and $w$ on the same path, then take such an edge $E_{1}$ with the condition that between $v$ and $w$ there are no pairs of vertices connected by edges in $M_{\Gamma}$. There is a vertex of another edge $E_{2}$ lying on the same path between $v$ and $w$, for otherwise there is a bigonal face, which is a contradiction. The edge $E_{2}$ lies in another connected component of $\partial Q\setminus\Gamma$. The other vertex of $E_{2}$ lies either on the same path, or on another path. In both cases it is clear from the octant representation that the circles are linked is the standard way (as in the Hopf link). This proves (1).
Now let $C_{\Gamma}$ consist of mutually unlinked circles. By (1) each edge of $M_{\Gamma}$ connects vertices on different paths of $\Gamma$. Let $\Gamma_{1}$, $\Gamma_{2}$ and $\Gamma_{3}$ be the paths of $\Gamma$ connecting the vertices $v$ and $w$. If there is an edge $E\in M_{\Gamma}$ with vertices $v_{i}$ and $v_{j}$ on $\Gamma_{i}$ and $\Gamma_{j}$ such that there are no vertices on these paths between $v$ and $v_{i}$ and $v$ and $v_{j}$, then $Q$ has a triangle incident to $v$ and if $Q\neq\Delta^{3}$, then $(Q,\Gamma)$ is obtained from some pair $(Q^{\prime},\Gamma^{\prime})$ by cutting off a vertex. The graph $G(Q^{\prime})$ is obtained from $G(Q)$ by deletion of $E$. The link $C_{\Gamma}$ is trivial if and only if $C_{\Gamma^{\prime}}$ is trivial. If $Q$ has no such edges $E$, then consider a vertex $v_{1}$ on $\Gamma_{1}$ closest to $v$. If this vertex is $w$, then all the edges of $M_{\Gamma}$ connect vertices on $\Gamma_{2}$ and $\Gamma_{3}$ and are “parallel”, in particular the first and the last edges are of the above type. A contradiction. Thus, $v_{1}$ belongs to some edge $E_{1}\in M_{\Gamma}$ with the other vertex $v_{2}$ lying on the other path, say $\Gamma_{2}$. By our assumption, there is a vertex $v_{3}$ between $v$ and $v_{2}$. Let $v_{3}$ be the closest vertex to $v$. Then $v_{3}\in E_{2}\in M_{\Gamma}$. The other vertex $v_{4}$ of $E_{2}$ belongs to $\Gamma_{3}$. Again by our assumption there is a vertex $v_{5}$ between $v$ and $v_{4}$, $v_{5}\in E_{3}\in M_{\Gamma}$. Let $v_{6}$ be the other vertex of $E_{3}$. Then $v_{6}\in\Gamma_{1}$ and $v_{1}$ lies between $v$ and $v_{6}$. The edges $E_{1}$, $E_{2}$ and $E_{3}$ correspond to Borromean rings in $C_{\Gamma}$. In particular $C_{\Gamma}$ is a nontrivial link. Thus, if $C_{\Gamma}$ does not contain Borromean rings, then $(Q,\Gamma)$ is obtained from $(\Delta^{3},\Gamma_{0})$ by a sequence of operations of cutting off a vertex. In particular, $C_{\Gamma}$ is trivial. Together with Example 7.11 this proves (2) and (3). ∎
Theorem 7.14.
Let $\Gamma$ be a Hamiltonian $K_{4}$-graph on a simple $3$-polytope $Q$. Then
  1. (1)
    the link $C_{\Gamma}$ consists of mutually unlinked circles if and only if $M_{\Gamma}$ splits into matchings $M_{\Gamma}(v_{i})$ corresponding to vertices of $K_{4}$, such that each matching consists of edges connecting the vertices on different paths of $K_{4}$ containing $v_{i}$ and for any two edges $E_{1}\in M_{\Gamma}(v_{i})$ and $E_{2}\in M_{\Gamma}(v_{j})$, $i\neq j$ the triangles $v_{i}*E_{1}$ and $v_{j}*E_{2}$ do not intersect;
  2. (2)
    if $C_{\Gamma}$ consists of mutually unlinked circles and is nontrivial, then it contains a triple of Borromean rings;
  3. (3)
    the link $C_{\Gamma}$ is trivial if and only if $(Q,\Gamma)$ is obtained from $(\Delta^{3},G(\Delta^{3}))$ by a sequence of operations of cutting off a vertex.
Proof.
If the condition of item (1) holds, then we can isotope all the matchings to be close to the corresponding vertices. Then near each vertex we have the theta-graph $\Gamma(v_{i})$ (obtained by shrinking to point the triangle of $\Gamma$ complementary to $v_{i}$) with the matching $M_{\Gamma}(v_{i})$, and the link $C_{\Gamma}$ consists of two copies of each link $C_{\Gamma}(v_{i})$ for all $i$ lying in disjoint disks. So $C_{\Gamma}$ consists of mutually unlinked circles.
Now let $C_{\Gamma}$ consist of mutually unlinked circles. If there is an edge in $M_{\Gamma}$ connecting the vertices on the same path of $\Gamma$, then consider such an edge $E_{2}$ with ends $w_{1}$ and $w_{2}$ and no other pairs of vertices between $w_{1}$ and $w_{2}$ connected by an edge in $M_{\Gamma}$. Since $Q$ has no bigons, there is a vertex $w_{3}$ between $w_{1}$ and $w_{2}$. This vertex is connected by an edge $E_{2}\in M_{\Gamma}$ to some other vertex $w_{4}$. It is clear from the representation of $K_{4}$ as the graph of the simplex $\Delta^{3}$ with vertices $v_{0}=(0,0,0)$, $v_{1}=(1,0,0)$, $v_{2}=(0,1,0)$, $v_{3}=(0,0,1)$ that either $E_{1}$ and $E_{2}$ correspond to $4$ pairs of circles linked in a standard way (if $w_{4}$ lies in the same path), or $E_{1}$ corresponds to $4$ and $E_{2}$ corresponds to $2$ unlinked circles such that each circle of the second type is linked in a standard way to two circles of the first type. A contradiction. Thus, each edge of $M_{\Gamma}$ connects two vertices on different paths. Then for each edge $E\in M_{\Gamma}$ there is a unique vertex $v_{i}$ of $K_{4}$ such that $E$ lies in the triangle $v_{i},v_{j},v_{k}]$ and connects the points on the paths $v_{i},v_{j}]$ and $v_{i},v_{k}]$. Denote this vertex $v(E)$. Now the proof of item (1) follows from
Lemma 7.15.
The link corresponding to two edges $E_{1},E_{2}\in M_{\Gamma}$ connecting vertices on different paths of $K_{4}$ is nontrivial if and only if $v(E_{1})\neq v(E_{2})$ and the triangles $v(E_{1})*E_{1}$ and $v(E_{2})*E_{2}$ intersect (equivalently, the segments between a vertex of $E_{1}$ and $v(E_{1})$ and a vertex of $E_{2}$ and $v(E_{2})$ lying both on $v(E_{1}),v(E_{2})]$ intersect). If this link is nontrivial, then it is the $4$-link chain like in Example 3.7.
Proof.
If $v(E_{1})=v(E_{2})$, then as in the above argument $C_{\Gamma}$ consists of two copies of the trivial link corresponding to two edges on the theta-graph. These copies lie in disjoint disks, so $C_{\Gamma}$ is trivial.
If $v(E_{1})\neq v(E_{2})$ and $v(E_{1})*E_{1}\cap v(E_{2})*E_{2}=\varnothing$, then $C_{\Gamma}$ is trivial, since it consists of four circles lying in disjoint balls.
If $v(E_{1})\neq v(E_{2})$ and $v(E_{1})*E_{1}\cap v(E_{2})*E_{2}\neq\varnothing$, then the edge $v(E_{1}),v(E_{2})]$ contains the vertex $w_{1}$ of $E_{1}$ and the vertex $w_{2}$ of $E_{2}$, and these vertices lie in the order $(v(E_{1}),w_{2},w_{1},v(E_{2}))$. Then each of the circles corresponding to $E_{1}$ is linked to each of the circles corresponding to $E_{2}$ in a standard way (like in the Hopf link) and these circles form the $4$-link chain. This finishes the proof. ∎
The proof of items (2) and (3) is the same as the proof of items (2) and (3) of Theorem 7.12. ∎
8. Eulerian theta-graphs and $K_{4}$-graphs on hyperbolic right-angled $3$-polytopes
Similarly to hyperbolic links corresponding to nonselfcrossing Eulerian cycles on ideal right-angled hyperbolic $3$-polytopes one can build hyperbolic links corresponding to nonselfcrossing Eulerian theta-graphs (or $K_{4}$-graphs) on right-angled hyperbolic polytopes $P$ of finite volume with $2$ (or $4$) finite vertices and all the other vertices ideal.
Definition 8.1.
An Eulerian theta-graph (or $K_{4}$-graph ) consists of $3$ (or $6$) paths connecting all pairs of finite vertices, and each edge of $P$ belongs exactly to one path. An Eulerian theta-graph (or $K_{4}$-graph) is nonselfcrossing if at each $4$-valent vertex is turns left or right.
Construction 8.2.
Let $\gamma$ be a nonselfcrossing Eulerian theta-graph (or $K_{4}$-graph) $\gamma$ on a $3$-polytope $P$ with $2$ (or $4$) 33-valent vertices and all the other vertices $4$-valent. Let us build a simple $3$-polytope $Q(P,\gamma)$ with a Hamiltonian theta-graph (or $K_{4}$-graph) $\Gamma_{\gamma}$ by the following rule. Substitute each $4$-valent vertex of the graph of $P$ by two vertices connected by an edge in such a way that each pair of successive edges of $\gamma$ at this vertex is incident to the same vertex of the new edge. The new graph satisfies the condition that each face is bounded by a simple edge-cycle, and if two boundary cycles of faces intersect, then by an edge. Thus by the Steinitz theorem this graph is a graph of a unique combinatorial simple polytope $Q(P,\gamma)$. Moreover, $\gamma$ corresponds to a Hamiltonian theta-graph (or $K_{4}$-graph) $\Gamma_{\gamma}$ on this polytope. The new edges form a matching in the graph of $Q$.
Theorem 8.3.
Let $P$ be a right-angled hyperbolic $3$-polytope of finite volume with $2$ (or $4$) finite vertices. Each nonselfcrossing Eulerian theta-graph (or $K_{4}$-graph) $\gamma$ corresponds to a link $C_{\gamma}\subset S^{3}$ such that its complement $S^{3}\setminus C_{\gamma}$ is homeomorphic to a hyperbolic manifold glued of $8$ (or $16$) copies of $P$. Moreover, $C_{\gamma}$ is the branch set of the $2$-fold branched covering $N(Q(P,\gamma),\widetilde{\Lambda}_{\Gamma_{\gamma}})\to S^{3}$.
Proof.
The proof is similar to the proof of Theorem 3.2. ∎
Problem 5.
Does any right-angled hyperbolic $3$-polytope of finite volume with $2$ (or $4$) finite vertices have a nonselfcrossing Eulerian theta-graph (or $K_{4}$-graph)? To enumerate all such graphs.
Problem 6.
To characterise Hamiltonian theta-graphs (or $K_{4}$-graphs) on simple $3$-polytopes corresponding to nonselfcrossing Eulerian theta-graphs (or $K_{4}$-graphs) on right-angled hyperbolic $3$-polytopes of finite volume.
9. Acknowledgements
The author is grateful to V.M. Buchstaber for stimulating questions that led to the expansion of the scope of this article. The author is grateful to N.V. Bogachev, A.A. Gaifullin, Vladimir Gorchakov, D.V. Gugnin, I.Yu. Limonchenko, A.D. Mednykh, T.E. Panov, M.V. Prasolov, D.V. Talalaev, D.A. Tsygankov, A.Yu. Vesnin for fruitful discussions.